### v1

@@ Abstract
Whole-slide images (WSIs) are critical for cancer diagnosis due to their ultra-high resolution and rich semantic content. However, their massive size and the limited availability of fine-grained annotations pose substantial challenges for conventional supervised learning. We propose DSAGL (Dual-Stream Attention-Guided Learning), a novel weakly supervised classification framework that combines a teacher-student architecture with a dual-stream design. DSAGL explicitly addresses instance-level ambiguity and bag-level semantic consistency by generating multi-scale attention-based pseudo labels and guiding instance-level learning. A shared lightweight encoder (VSSMamba) enables efficient long-range dependency modeling, while a fusion-attentive module (FASA) enhances focus on sparse but diagnostically relevant regions. We further introduce a hybrid loss to enforce mutual consistency between the two streams. Experiments on CIFAR-10, NCT-CRC, and TCGA-Lung datasets demonstrate that DSAGL consistently outperforms state-of-the-art MIL baselines, achieving superior discriminative performance and robustness under weak supervision.
@@ 1 Introduction
In the field of digital medicine, ultra-high-resolution whole slide images (WSIs) have become a core imaging modality in modern pathological diagnosis due to their comprehensive and information-rich representation [35, 42, 8, 50]. WSIs can capture microscopic tissue structures with exceptional precision, providing critical support for cancer diagnosis and treatment planning. However, their massive size imposes substantial computational burdens [11, 43]. The complexity of tissue structures further complicates the accurate identification of abnormal regions, which often exhibit diffuse and scattered morphological features that hinder automated analysis. Traditional supervised learning methods rely heavily on exhaustive annotations by expert pathologists—a process that is not only time-consuming and costly but also subject to significant inter-observer variability, ultimately compromising model generalization. To address these challenges, weakly supervised learning approaches have gained increasing attention in computational pathology [27, 29, 34, 38, 44]. By leveraging easily accessible patient-level labels, these methods reduce the dependency on high-quality annotations and offer a promising pathway toward efficient and accurate analysis of large-scale pathological images.
In recent years, significant progress has been made in whole-slide image (WSI) classification for computational pathology, achieving notable progress. Among these approaches, multiple instance learning (MIL) [18, 19, 26, 31] has emerged as a dominant paradigm. MIL-based methods typically regard a WSI as a “bag” composed of numerous image patches and leverage patient-level labels to guide instance-level learning [51, 3]. This formulation not only makes effective use of coarse-grained annotations but also aligns well with clinical diagnostic practices, significantly reducing the cost of data labeling while delivering promising results across various tasks. However, existing MIL methods often rely on the assumption that instances within a bag are independent and identically distributed (i.i.d.), or adopt simplistic aggregation strategies such as attention-based pooling [19]. These approaches are limited in their ability to capture complex spatial dependencies and contextual information that are critical for accurately identifying diagnostic regions. The lack of long-range modeling across spatially distant patches hinders precise localization and recognition of relevant lesions, especially in real-world scenarios characterized by sparse and heterogeneous pathological patterns. Moreover, a fundamental trade-off persists between instance-level discriminability and bag-level classification performance, which compromises the model’s generalization and clinical applicability. Therefore, under weak supervision, designing more expressive and context-aware architectures that can jointly enhance instance-level and bag-level performance remains a central challenge in computational pathology.
@@
To address the aforementioned challenges, we propose DSAGL (Dual-Stream Attention-Guided Learning), a novel weakly supervised classification framework based on a dual-stream, teacher–student architecture with attention guidance. DSAGL is designed to capture inter-instance correlations while jointly enhancing both instance-level discriminability and bag-level classification performance in MIL models. The framework consists of a unified feature extraction backbone and a dual-stream structure. Specifically, the teacher stream employs a multi-scale attention mechanism to focus on diagnostically relevant instances, enabling accurate bag-level classification while generating high-quality soft pseudo labels. Guided by these pseudo labels, the student stream performs instance-level classification, thereby improving fine-grained discriminative capability. Moreover, as shown in Fig. 1(b), DSAGL employs an alternating training strategy between the teacher and student networks, in contrast to the sequential paradigm adopted by conventional dual-stream models. This collaborative optimization enhances semantic consistency and further improves the overall model performance.
@@
  • We propose a novel weakly supervised classification framework, DSAGL, which integrates a dual-stream architecture with a teacher–student mechanism to jointly optimize instance-level discriminability and bag-level classification performance, addressing the limitations of conventional MIL methods in semantic modeling.
@@ 2.1 Multiple Instance Learning for WSI Classification
Multiple instance learning (MIL) [47, 40, 40] has become a widely adopted weakly supervised paradigm in computational pathology, particularly well-suited for whole slide images (WSIs), which are characterized by extremely high resolution and large-scale tissue content. In MIL, each WSI is divided into a set of patches (instances) grouped into a bag, with supervision provided only at the bag level. This eliminates the need for dense annotations and significantly reduces manual labeling costs.
@@ 2.2 Dual-Stream Architectures and Knowledge Distillation
Prior to the emergence of dual-stream architectures, extensive research explored the integration of knowledge distillation (KD) into single-branch or loosely coupled frameworks to enhance the representational capacity and training stability of student models. For example, Co-MIL [20] employs a multi-teacher collaborative distillation strategy, where separate teacher models extract global and local semantic information to guide the student through cross-model supervision. GLMC [10] introduces bidirectional guidance between global and local branches, thereby improving the model’s ability to capture multi-scale lesion features. Self-Distillation utilizes the historical states of the student model as an implicit teacher, enabling knowledge transfer without additional parameter overhead—an approach particularly suitable for resource-constrained settings. ReNA-KD [52] incorporates region-structured attention maps into the distillation process, enhancing spatial awareness and improving recognition of heterogeneous lesion patterns.
@@ 2.3 Long-Range Dependency Modeling in WSI
Whole Slide Images (WSIs) are composed of thousands of image patches and exhibit strong non-local structural characteristics. In pathological diagnosis, spatially distant regions often present semantic dependencies—for example, between tumor boundaries and cores, or between inflamed tissues and surrounding microenvironments. Modeling such long-range dependencies is critical for improving classification accuracy.
@@
Derived from the Mamba framework, several variants have been developed. Vision Mamba [53] integrates convolutional pre-processing and spatial shift mechanisms to enhance spatial awareness. T-Mamba [17] introduces temporal decoupling for efficient modeling of non-uniform sequential data. U-Mamba [32] extends the architecture for biomedical image segmentation by incorporating a U-shaped encoder–decoder design with Mamba blocks, enabling long-range context modeling while preserving spatial resolution.
@@ 3.1.1 Multiple Instance Learning (MIL)
Each Whole Slide Image (WSI) is divided into a bag of image patches, where each patch is treated as an instance [19, 27, 29, 41]. A bag is defined as $B_{i}=\{x_{i,1},\dots,x_{i,N_{i}}\}$, where $x_{i,j}\in\mathbb{R}^{\texttt{d}}$ denotes the d-dimensional feature vector of the $j$-th instance in bag $i$, and $N_{i}$ is the number of instances.
@@ 3.2.1 DSAGL architecture
To achieve efficient and interpretable pathological image classification, we propose a Dual-Stream Attention-Guided Learning Weakly Supervised Framework (DSAGL). As illustrated in Fig. 3, the DSAGL architecture consists of two main components: a feature extraction module and a dual-stream network comprising a teacher branch and a student branch. Specifically, we first employ the Vision-aware Skip-connected Selective Mamba encoder (VSSMamba) to extract preliminary instance-level features from the input images. These features are then processed independently by the teacher and student branches. In the teacher branch, we introduce a Fusion-Attentive Scale-Aware module (FASA) to generate bag-level predictions and compute attention scores, which are subsequently used to construct high-quality instance-level pseudo-labels. The student branch focuses on instance-level classification, using the pseudo-labels from the teacher as supervision signals. By alternately updating the teacher and student networks, DSAGL facilitates efficient and collaborative feature sharing and representation learning.
@@ 3.2.2 Alternating Training Strategy
To enhance the collaborative optimization between the teacher and student networks under weak supervision, DSAGL adopts an alternating training strategy. Rather than employing conventional joint or sequential training paradigms, this strategy updates the two branches in an interleaved manner, enabling iterative refinement of pseudo labels and feature representations throughout the training process. During training, the teacher network is updated at fixed intervals, while the student network is optimized continuously using the pseudo labels generated by the teacher. This staged update pattern mitigates potential gradient interference between the two branches and contributes to stable convergence. Moreover, by decoupling the training dynamics, each network can specialize in its respective task: global bag-level prediction for the teacher and fine-grained instance-level discrimination for the student.
This alternating strategy offers several key advantages. First, it facilitates dynamic refinement of pseudo labels, which are progressively improved throughout training, thereby reducing the impact of noise in the early stages. Second, it enables a clear decoupling of tasks between the two branches, allowing the teacher to focus on holistic aggregation while the student learns detailed discriminative features. Third, the staggered updates improve overall training stability by avoiding simultaneous gradient conflicts. Finally, this design proves particularly effective in weakly supervised scenarios, where sparse or noisy labels can severely degrade performance.
@@ 3.4.3 Student Network
To guide the student in identifying diagnostically relevant regions, the attention score $a_{i,j}$ provided by the teacher is passed through a normalization module, NormProb, which converts it into a probabilistic pseudo label $\hat{z}_{i,j}$. This process stabilizes the supervision signal and helps mitigate early-stage noise in the attention distribution. The training of the student network is supervised using a hybrid loss function $\mathcal{L}_{Hybrid}$, which combines cross-entropy loss [33] with Kullback–Leibler divergence [24] in a weighted formulation. This design allows the network to benefit from both hard pseudo labels and soft distribution alignment. The resulting student loss is denoted by $\mathcal{L}_{S}$, and is computed as follows:
@@ 3.4.4 Hybrid Loss Function
The student cross-entropy loss is defined as:
$\mathcal{L}_{CE,student}=-\frac{1}{N}\sum_{i,j}\left[w_{n}(1-\hat{z}_{i,j})\log(p_{i,j,0}+\epsilon)+(1-w_{n})\hat{z}_{i,j}\log(p_{i,j,1}+\epsilon)\right],$ (24)
@@ 4.1.1 Cifar-10
Cifar-10 [23] is a standard image classification dataset consisting of 60,000 color images of size $32\times 32$ pixels, divided into 10 classes, with 6,000 images per class. Among them, 50,000 images are used for training and 10,000 for testing. The training set is organized into five batches of 10,000 images each, while the test set constitutes a separate batch. In weakly supervised learning tasks, Cifar-10 is commonly utilized to simulate whole-slide images (WSIs) by randomly assembling images from different classes to construct synthetic datasets. In our setting, we simulate WSIs by randomly combining images from different classes within the dataset.
@@ 4.1.3 TCGA-Lung
The TCGA (The Cancer Genome Atlas) lung cancer dataset [26] is an important publicly available resource for histopathological image analysis, encompassing two major subtypes of non-small cell lung cancer: lung adenocarcinoma (LUAD) and lung squamous cell carcinoma (LUSC). The dataset comprises a total of 1,054 high-resolution whole-slide images (WSIs) stained with hematoxylin and eosin (H&E), provided in the svs format and accessible through the data portal of the National Cancer Institute of the United States. At 20×20$\times$20 × magnification, each WSI is partitioned into approximately 5.2 million image patches of size $224\times 224$ pixels. Each WSI constitutes a bag in the multiple instance learning setting, containing a varying number of tumor and normal patches.

### v2

@@ Abstract
Whole Slide Images are critical for cancer diagnosis due to their ultra-high resolution and rich semantic content. However, their massive size and the limited availability of fine-grained annotations pose substantial challenges for conventional supervised learning. We propose the dual-stream attention-guided learning framework called DSAGL, a novel weakly supervised classification framework that combines a teacher-student architecture with a dual-stream design. DSAGL explicitly addresses instance-level ambiguity and bag-level semantic consistency by generating multi-scale attention-based pseudo labels and guiding instance-level learning. A shared lightweight encoder enables efficient long-range dependency modeling, while a fusion-attentive module enhances focus on sparse but diagnostically relevant regions. We further introduce a hybrid loss to enforce mutual consistency between the two streams. Experiments on CIFAR-10, NCT-CRC, and TCGA-Lung datasets demonstrate that DSAGL consistently outperforms state-of-the-art MIL baselines, achieving superior discriminative performance and robustness under weak supervision.
@@ 1 Introduction
In the field of digital medicine, ultra-high-resolution Whole Slide Images (WSIs) have become a core imaging modality in modern pathological diagnosis due to their comprehensive and information-rich representation [35, 42, 8, 49]. WSIs can capture microscopic tissue structures with exceptional precision, providing critical support for cancer diagnosis and treatment planning. However, their massive size imposes substantial computational burdens [11, 43]. The complexity of tissue structures further complicates the accurate identification of abnormal regions, which often exhibit diffuse and scattered morphological features that hinder automated analysis. Traditional supervised learning methods rely heavily on exhaustive annotations by expert pathologists-a process that is not only time-consuming and costly but also subject to significant inter-observer variability, ultimately compromising model generalization. To address these challenges, weakly supervised learning approaches have gained increasing attention in computational pathology [27, 29, 34, 38, 44]. By leveraging easily accessible patient-level labels, these methods reduce the dependency on high-quality annotations and offer a promising pathway toward efficient and accurate analysis of large-scale pathological images.
In recent years, significant progress has been made in WSI classification for computational pathology, achieving notable progress. Among these approaches, multiple instance learning (MIL) [18, 19, 26, 31] has emerged as a dominant paradigm. MIL-based methods typically regard a WSI as a "bag" composed of numerous image patches and leverage patient-level labels to guide instance-level learning [50, 3]. This formulation not only makes effective use of coarse-grained annotations but also aligns well with clinical diagnostic practices, significantly reducing the cost of data labeling while delivering promising results across various tasks. However, existing MIL methods often rely on the assumption that instances within a bag are independent and identically distributed (i.i.d.), or adopt simplistic aggregation strategies such as attention-based pooling [19]. These approaches are limited in their ability to capture complex spatial dependencies and contextual information that are critical for accurately identifying diagnostic regions. The lack of long-range modeling across spatially distant patches hinders precise localization and recognition of relevant lesions, especially in real-world scenarios characterized by sparse and heterogeneous pathological patterns. Moreover, a fundamental trade-off persists between instance-level discriminability and bag-level classification performance, which compromises the model’s generalization and clinical applicability. Therefore, under weak supervision, designing more expressive and context-aware architectures that can jointly enhance instance-level and bag-level performance remains a central challenge in computational pathology.
@@
To address the aforementioned challenges, we propose DSAGL (Dual-Stream Attention-Guided Learning), a novel weakly supervised classification framework based on a dual-stream, teacher-student architecture with attention guidance. DSAGL is designed to capture inter-instance correlations while jointly enhancing both instance-level discriminability and bag-level classification performance in MIL models. The framework consists of a unified feature extraction backbone and a dual-stream structure. Specifically, the teacher stream employs a multi-scale attention mechanism to focus on diagnostically relevant instances, enabling accurate bag-level classification while generating high-quality soft pseudo labels. Guided by these pseudo labels, the student stream performs instance-level classification, thereby improving fine-grained discriminative capability. Moreover, as shown in Fig. 1(b), DSAGL employs an alternating training strategy between the teacher and student networks, in contrast to the sequential paradigm adopted by conventional dual-stream models. This collaborative optimization enhances semantic consistency and further improves the overall model performance.
@@
  • We propose a novel weakly supervised classification framework, DSAGL, which integrates a dual-stream architecture with a teacher-student mechanism to jointly optimize instance-level discriminability and bag-level classification performance, addressing the limitations of conventional MIL methods in semantic modeling.
@@ 2.1 Multiple Instance Learning for WSI Classification
Multiple instance learning (MIL) [47, 40, 40] has become a widely adopted weakly supervised paradigm in computational pathology, particularly well-suited for WSIs, which are characterized by extremely high resolution and large-scale tissue content. In MIL, each WSI is divided into a set of patches (instances) grouped into a bag, with supervision provided only at the bag level. This eliminates the need for dense annotations and significantly reduces manual labeling costs.
@@ 2.2 Dual-Stream Architectures and Knowledge Distillation
Prior to the emergence of dual-stream architectures, extensive research explored the integration of knowledge distillation (KD) into single-branch or loosely coupled frameworks to enhance the representational capacity and training stability of student models. For example, Co-MIL [20] employs a multi-teacher collaborative distillation strategy, where separate teacher models extract global and local semantic information to guide the student through cross-model supervision. GLMC [10] introduces bidirectional guidance between global and local branches, thereby improving the model’s ability to capture multi-scale lesion features. Self-Distillation utilizes the historical states of the student model as an implicit teacher, enabling knowledge transfer without additional parameter overhead-an approach particularly suitable for resource-constrained settings. ReNA-KD [51] incorporates region-structured attention maps into the distillation process, enhancing spatial awareness and improving recognition of heterogeneous lesion patterns.
@@ 2.3 Long-Range Dependency Modeling in WSI
WSIs are composed of thousands of image patches and exhibit strong non-local structural characteristics. In pathological diagnosis, spatially distant regions often present semantic dependencies-for example, between tumor boundaries and cores, or between inflamed tissues and surrounding microenvironments. Modeling such long-range dependencies is critical for improving classification accuracy.
@@
Derived from the Mamba framework, several variants have been developed. Vision Mamba [52] integrates convolutional pre-processing and spatial shift mechanisms to enhance spatial awareness. T-Mamba [17] introduces temporal decoupling for efficient modeling of non-uniform sequential data. U-Mamba [32] extends the architecture for biomedical image segmentation by incorporating a U-shaped encoder-decoder design with Mamba blocks, enabling long-range context modeling while preserving spatial resolution.
@@ 3.1.1 Multiple Instance Learning (MIL)
Each WSI is divided into a bag of image patches, where each patch is treated as an instance [19, 27, 29, 41]. A bag is defined as $B_{i}=\{x_{i,1},\dots,x_{i,N_{i}}\}$, where $x_{i,j}\in\mathbb{R}^{\texttt{d}}$ denotes the d-dimensional feature vector of the $j$-th instance in bag $i$, and $N_{i}$ is the number of instances.
@@ 3.2.1 DSAGL architecture
To enable efficient and interpretable pathological image classification, we propose a Dual-Stream Attention-Guided Learning framework (DSAGL) for weakly supervised learning. As illustrated in Fig. 3, the DSAGL architecture comprises two main components: a feature extraction module and a dual-stream network consisting of a teacher branch and a student branch. Specifically, we first adopt the Vision-aware Skip-connected Selective Mamba encoder (VSSMamba) to extract initial instance-level features from the input patches. These features are then independently processed by the teacher and student branches. In the teacher branch, we introduce the Fusion-Attentive Scale-Aware (FASA) module to generate bag-level predictions and compute attention scores, which are subsequently used to construct high-quality instance-level pseudo labels. The student branch focuses on fine-grained instance-level classification, guided by the pseudo labels provided by the teacher. Through an alternating training strategy between the two branches, DSAGL enables efficient collaborative feature learning and semantic alignment under weak supervision.
@@ 3.2.2 Alternating Training Strategy
To enhance collaborative optimization between the teacher and student networks under weak supervision, DSAGL adopts an alternating training strategy. Instead of relying on conventional joint or sequential training paradigms, this approach updates the two branches in an interleaved manner, enabling iterative refinement of both pseudo labels and feature representations throughout the training process. During training, the teacher network is updated at fixed intervals, while the student network is continuously optimized using the pseudo labels generated by the teacher. This staged update pattern mitigates potential gradient interference between the two branches and contributes to more stable convergence. By decoupling the training dynamics, each network is able to specialize in its respective task: global bag-level prediction for the teacher and fine-grained instance-level discrimination for the student.
This alternating strategy offers several key advantages. First, it facilitates dynamic refinement of pseudo labels, which are progressively improved during training, thereby reducing the impact of early-stage noise. Second, it enables clear task decoupling, allowing the teacher to focus on holistic feature aggregation while the student learns detailed discriminative representations. Third, the staggered updates improve overall training stability by preventing simultaneous gradient conflicts. Finally, this design proves particularly effective in weakly supervised scenarios, where sparse or noisy labels can severely compromise learning performance.
@@ 3.4.3 Student Network
To guide the student in identifying diagnostically relevant regions, the attention score $a_{i,j}$ provided by the teacher is passed through a normalization module, NormProb, which converts it into a probabilistic pseudo label $\hat{z}_{i,j}$. This process stabilizes the supervision signal and helps mitigate early-stage noise in the attention distribution. The training of the student network is supervised using a hybrid loss function $\mathcal{L}_{Hybrid}$, which combines cross-entropy loss [33] with Kullback-Leibler divergence [24] in a weighted formulation. This design allows the network to benefit from both hard pseudo labels and soft distribution alignment. The resulting student loss is denoted by $\mathcal{L}_{S}$, and is computed as follows:
@@ 3.4.4 Hybrid Loss Function
$\begin{split}\mathcal{L}_{CE,student}=\;&-\frac{1}{N}\sum_{i,j}\Big{[}w_{n}(1-\hat{z}_{i,j})\log(p_{i,j,0}+\epsilon)\\ &\qquad+\,(1-w_{n})\hat{z}_{i,j}\log(p_{i,j,1}+\epsilon)\Big{]}\end{split}$ (24)
@@ 4.1.1 Cifar-10
Cifar-10 [23] is a standard image classification dataset consisting of 60,000 color images of size $32\times 32$ pixels, divided into 10 classes, with 6,000 images per class. Among them, 50,000 images are used for training and 10,000 for testing. The training set is organized into five batches of 10,000 images each, while the test set constitutes a separate batch. In weakly supervised learning tasks, Cifar-10 is commonly utilized to simulate WSIs by randomly assembling images from different classes to construct synthetic datasets. In our setting, we simulate WSIs by randomly combining images from different classes within the dataset.
@@ 4.1.3 TCGA-Lung
The TCGA (The Cancer Genome Atlas) lung cancer dataset [26] is an important publicly available resource for histopathological image analysis, encompassing two major subtypes of non-small cell lung cancer: lung adenocarcinoma (LUAD) and lung squamous cell carcinoma (LUSC). The dataset comprises a total of 1,054 high-resolution WSIs stained with hematoxylin and eosin (H&E), provided in the svs format and accessible through the data portal of the National Cancer Institute of the United States. At 20×20$\times$20 × magnification, each WSI is partitioned into approximately 5.2 million image patches of size $224\times 224$ pixels. Each WSI constitutes a bag in the multiple instance learning setting, containing a varying number of tumor and normal patches.
@@ 4.2 Experimental Details
4.3 Comparison Methods
To systematically evaluate the effectiveness of the proposed DSAGL framework, we compare it against a range of representative multiple instance learning (MIL) approaches. Based on their architectural characteristics, these methods can be categorized into three groups:
Classical MIL Models. Representative methods such as MI-Net [48] and RNN-MIL [4] adopt fixed pooling strategies (e.g., max or mean pooling) to aggregate instance features into bag-level representations. While computationally efficient, these methods lack mechanisms for modeling discriminative regions explicitly and often underperform in weakly supervised scenarios.
Attention-based MIL Models. Attention-MIL and its variants, including ABMIL and Gated-ABMIL [19], introduce learnable attention modules to assign instance-specific weights, enabling flexible aggregation and improved interpretability. CLAM [31] further enhances localization by incorporating attention pruning to focus on diagnostically relevant patches.
Transformer-based or Advanced MIL Models. This category enhances MIL performance by integrating global context modeling and architectural refinements. TransMIL [39] leverages self-attention to capture inter-instance dependencies; IBMIL [29] introduces region-level interventions for improved instance discrimination; R2T-MIL [45] adopts token-level guidance to promote region-aware learning. Our proposed DSAGL also falls into this category, combining dual-stream optimization, attention-guided pseudo supervision, and efficient long-range modeling via the VSSMamba encoder.
All models are evaluated on the CIFAR-10 and NCT-CRC datasets under varying positive instance ratios. Specifically, CIFAR-10 is tested under ratios ranging from 1% to 70%, while NCT-CRC uses ratios from 10% to 70%. For the TCGA-Lung dataset, which lacks instance-level annotations, only bag-level classification performance is reported. Baseline results are either reproduced under our unified experimental settings or cited directly from existing studies on weakly supervised WSI classification [36].
4.4 Experimental Results
4.4.1 Results on the CIFAR-10 Dataset
To evaluate model performance under controllable weak supervision, we conduct extensive experiments on the synthetic CIFAR-10 dataset with varying positive patch ratios (1% to 70%). As summarized in Table 1, DSAGL consistently achieves superior classification performance across all settings. Specifically, at 10%, 20%, 50%, and 70% positive ratios, DSAGL attains instance-level AUCs of 0.9037, 0.9287, 0.9092, and 0.9103, respectively-outperforming strong baselines such as ABMIL, CLAM, and advanced Transformer-based models including TransMIL and IBMIL. At the bag level, DSAGL achieves perfect or near-perfect classification with AUC = 1.0000 in the 20%, 50%, and 70% settings, demonstrating remarkable robustness against label imbalance and noisy instances. Fig. 5 further visualizes the comparative results across positive ratios, validating the consistent performance advantage of DSAGL over existing methods.
To analyze the learning dynamics, we plot the test AUC trajectories over 300 training epochs under the 20% positive patch ratio setting (Fig. 6). All baseline models are reimplemented using publicly available code under identical settings, including the same feature extractor, optimizer, and training schedule, to ensure fair comparison. DSAGL demonstrates fast convergence and stable improvement throughout training, achieving consistently high AUCs at both the instance and bag levels. In contrast, classical methods such as MI-Net, ABMIL, and Gated-ABMIL converge slowly and exhibit noticeable oscillations, while more advanced models like CLAM, TransMIL, and IBMIL display erratic fluctuations or early performance stagnation despite achieving moderate peak AUCs.
These observations highlight the training stability and robustness of DSAGL, attributed to its dual-stream architecture and attention-guided pseudo supervision. The teacher stream, equipped with the FASA module, produces structured multi-scale soft attention maps as reliable supervision signals, while the student stream benefits from fine-grained learning initialized by attention priors. Moreover, the alternating update strategy enables progressive refinement of pseudo labels, mitigating the effect of early noise and enhancing label reliability. This synergy contributes to more stable optimization under weak supervision and improved generalization on noisy instance distributions.